OpenStudy (anonymous):

For all nonzero r ,t, and z values, 16r^3tz^5/-4rt^3z^2 is equal to what?

OpenStudy (anonymous):

\[\frac{ 16r ^{3}tz ^{5} }{ -4rt ^{3}z ^{2} }\]

OpenStudy (anonymous):

Do this:\[ \frac{t}{t^3} = t^{1-3} = t^{-2} = \frac{1}{t^2} \]For all the variables.

OpenStudy (anonymous):

For the non-variables, prime factorization helps: \[ 16 = 2^4, -4 = -2^2 \]

OpenStudy (anonymous):

\[ \frac{16}{-4}=\frac{2^4}{-2^2} = -2^{4-2} = -2^2 = -4 \]

OpenStudy (anonymous):

Well, it's pretty obvious that \(16/-4=-4\), but the point is to show the pattern here.

OpenStudy (anonymous):

So then it would be \[\frac{ -4r ^{2}z ^{3} }{ t ^{2} }\]

OpenStudy (anonymous):

that looks right.

OpenStudy (anonymous):

Thank you very much for the help!